The height h (in feet) of an object that is dropped from the height of s feet is given by the formula h = s - 16t 2 , where t is the time the object has been falling. A 5 foot tall woman on a sidewalk looks directly overhead and sees a window washer drop a bottle from the 3 story. How long does she have to get out of the way? Round to the nearest tenth. (A story is 12 feet.)
Answer by stanbon(48546)

The height h (in feet) of an object that is dropped from the height of s feet is given by the formula h = s - 16t 2 , where t is the time the object has been falling. A 5 foot tall woman on a sidewalk looks directly overhead and sees a window washer drop a bottle from the 3 story. How long does she have to get out of the way? Round to the nearest tenth. (A story is 12 feet.)
--------------------
h(t) = s - 16t^2
------------------
5 = 36 -16t^2
-16t^2 = -31
t^2 = 31/16
t = (sqrt(31))/4
t = 1.39.. seconds
========================
Cheers,
Stan H.